# DeblurSDI: Blind Image Deblurring Using Self-diffusion

### v1

@@ Abstract
Blind image deconvolution is a challenging ill-posed inverse problem, where both the latent sharp image and the blur kernel are unknown. Traditional methods often rely on handcrafted priors, while modern deep learning approaches typically require extensive pre-training on large external datasets, limiting their adaptability to real-world scenarios. In this work, we propose DeblurSDI, a zero-shot, self-supervised framework based on self-diffusion (SDI) that requires no prior training. DeblurSDI formulates blind deconvolution as an iterative reverse self-diffusion process that starts from pure noise and progressively refines the solution. At each step, two randomly-initialized neural networks are optimized continuously to refine the sharp image and the blur kernel. The optimization is guided by an objective function combining data consistency with a sparsity-promoting $\ell_{1}$-norm for the kernel. A key innovation is our noise scheduling mechanism, which stabilizes the optimization and provides remarkable robustness to variations in blur kernel size. These allow DeblurSDI to dynamically learn an instance-specific prior tailored to the input image. Extensive experiments demonstrate that DeblurSDI consistently achieves superior performance, recovering sharp images and accurate kernels even in highly degraded scenarios.
@@ 1 Introduction
Image deblurring is a fundamental task in computer vision, aiming to recover a sharp image from a blurred observation. The blur can arise from various sources, such as camera shake, object motion, or defocus, whose degradation model is typically formulated with the convolution between the latent sharp image and the blur kernel. In blind image deconvolution, both image and kernel are unknown and must be jointly estimated from a blurry image alone, which is a particularly challenging ill-pose inverse problem. Traditional blind deblurring methods rely on handcrafted priors, such as gradient sparsity (Fergus et al., 2006) or variational Bayesian frameworks (Levin et al., 2007), and are effective for simple blur scenarios but can be sensitive to complex blurs and noise. In contrast, modern deep learning approaches leverage large-scale datasets to learn complex mappings from blurred to sharp images (Nah et al., 2017; Kupyn et al., 2018; Laroche et al., 2024). However, these methods often require extensive pre-training, which limits their generalization to new blur types or real-world conditions with varying blurring conditions. More recently, a different line of research has emerged that applies the principles of Deep Image Prior (DIP) to this task (Ren et al., 2020). While these methods optimize a randomly-initialized network to fit a single observed image, leveraging its implicit structure as a regularizer without external training data, they are often sensitive to the joint optimization process and require careful adjustment of parameters, such as the kernel size, which poses a significant challenge for practical applications.
In this work, we introduce DeblurSDI, a novel framework that uniquely addresses the challenges of blind image deblurring. Our primary contribution is the reformulation of this ill-posed problem as a zero-shot, reverse self-diffusion process, which allows for dynamic, instance-specific optimization without reliance on large training datasets. To overcome the inherent instability of jointly estimating the image and blur kernel, we introduce a novel noise scheduler for both the image and kernel prior networks. This design is central to our method’s success, providing exceptional robustness to variations in blur kernel size—a significant limitation of prior methods. Through extensive experiments, we demonstrate that DeblurSDI consistently recovers sharp images and accurate kernels.
@@ 2 Related Work
Blind image deblurring has been extensively studied under probabilistic frameworks, with Variational Bayes (VB) methods being a prominent approach (Fergus et al., 2006). VB-based techniques model both the latent sharp image and the blur kernel as random variables and estimate their posterior distributions, often incorporating priors for image gradients. A drawback is that they can be more sensitive to initialization and may converge to local minima, limiting their performance in challenging scenario. Deep Image Prior (DIP) methods have also been explored, where a randomly-initialized convolutional network is optimized to fit the blurred image, using the network’s implicit structure as a regularizer (Ulyanov et al., 2018). These methods can be prone to overfitting to noise and artifacts if not properly regularized or early-stopped.
In recent years, deep learning techniques have emerged as powerful alternatives to traditional VB and MAP-based blind deblurring methods. Unlike classical approaches that rely on hand-crafted priors, deep networks are capable of learning complex image statistics and blur distributions directly from large-scale datasets. Early works, such as multi-scale convolutional networks proposed by  Nah et al. (2017), aim to directly restore sharp images from blurry inputs in an end-to-end fashion, bypassing explicit kernel estimation. These methods leverage hierarchical features to handle spatially varying blur and have demonstrated significant improvements over conventional approaches in both qualitative and quantitative evaluations. Another line of research focuses on explicitly predicting the blur kernel using neural networks. Sun et al. (2015) proposed a patch-wise kernel prediction strategy, where the network estimates local blur kernels that are subsequently used for non-blind deconvolution. More recent works explore unsupervised or self-supervised learning schemes, incorporating cycle-consistency or reconstruction losses to mitigate domain gaps between synthetic and real-world data Kupyn et al. (2018); Tao et al. (2018).
Recently, diffusion models have emerged as powerful generative priors for solving blind image deblurring. (Murata et al., 2023) proposed GibbsDDRM, a partially collapsed Gibbs sampler that alternates between diffusion-based image restoration and analytical updates of the degradation parameters. (Whang et al., 2022) introduced a stochastic refinement strategy that leverages diffusion sampling to iteratively improve deblurring quality. (Ren et al., 2023) further extended this line of work with a multiscale, structure-guided diffusion model that improves convergence and preserves fine details. More recently, (Laroche et al., 2024) combined diffusion models with an Expectation–Maximization framework, where the E-step samples the latent sharp image distribution using a diffusion prior and the M-step updates the blur kernel via a MAP estimation. Overall, deep learning approaches have demonstrated remarkable performance in blind image deblurring by learning flexible priors for both latent images and blur kernels. However, their performance can still degrade when faced with severe blur, unseen motion patterns, or domain shifts between training and test data.
3 Blind deblurring via self-diffusion
@@
This work addresses blind image deblurring by extending the self-diffusion framework, a general, zero-shot approach for solving inverse problems Luo & Huang (2025). The original self-diffusion method was designed for non-blind problems where the degradation operator is known. Our key contribution is to adapt this powerful framework to the more challenging blind deconvolution setting, where the degradation operator—specifically the blur kernel—is also unknown. To achieve this, we introduce a novel process that jointly recovers the clean image and the blur kernel by optimizing two coupled, untrained neural networks within a self-contained reverse diffusion process.
3.1 Self-Diffusion
@@
Self-diffusion is a training-free paradigm designed to solve general linear inverse problems of the form $\mathcal{A}\mathbf{x}_{\text{true}}=\mathbf{y}$, where $\mathcal{A}$ is a known forward operator, $\mathbf{x}_{\text{true}}$ is the unknown solution, and $\mathbf{y}$ is the observation. It operates via an iterative reverse diffusion process that starts from pure Gaussian noise. At each step $t$, a noisy version of the current estimate $\mathbf{x}_{t}$ is created with
$\hat{\mathbf{x}}_{t}=\mathbf{x}_{t}+\sigma_{t}\cdot\epsilon_{t}~.$
@@
A single, randomly initialized network at the first time step—the denoiser $D_{\theta}$—is then optimized continuously by minimizing a data fidelity loss with respect to the original observation $\mathbf{y}$ for each time step $t$,
$\mathcal{L}_{t}(\theta)=\|\mathcal{A}D_{\theta,t}(\hat{\mathbf{x}}_{t})-\mathbf{y}\|_{2}^{2}~.$
@@
3.2 Joint Image and Kernel Estimation
The forward model for blind image deblurring is given by
$\mathbf{y}=\mathbf{x}_{\text{true}}\circledast\mathbf{k}+n~,$
where the sharp image $\mathbf{x}_{\text{true}}\in\mathbb{R}^{H\times W\times C}$ and the blur kernel $\mathbf{k}\in\mathbb{R}^{K\times K}$ are both unknown. To adapt the self-diffusion framework to this blind setting, we must estimate both variables simultaneously. We achieve this by employing two dedicated, randomly initialized networks: an image denoiser $D_{\theta}$ to restore $\mathbf{x}_{\text{true}}$, and a kernel generator $G_{\phi}$ to produce $\mathbf{k}$. Our method simulates a reverse diffusion process over $T$ discrete time steps, starting with random noise for both the image estimate, $\mathbf{x}_{T}$, and the kernel estimate, $\mathbf{z}_{T}$. At each time step $t\in\{T,T-1,...,1\}$, the current estimates are perturbed with scheduled noise,
$\hat{\mathbf{x}}_{t}=\mathbf{x}_{t}+\sigma_{t}\cdot\epsilon_{x},\quad\text{and}\quad\hat{\mathbf{z}}_{t}=\mathbf{z}_{t}+\sigma^{\prime}_{t}\cdot\epsilon_{z},$
where $\epsilon_{x}\sim\mathcal{N}(0,\mathbf{I})$ and $\epsilon_{z}\sim\mathcal{N}(0,\mathbf{I})$. The noise schedule is $\sigma_{t}=\sqrt{1-\bar{\alpha}_{t}},~\text{where}~\bar{\alpha}_{t}=\mathop{\textstyle\prod}\nolimits_{i=0}^{t}(1-\beta_{i})~\text{and}~\beta_{t}=\beta_{\text{end}}+\frac{t}{T-1}(\beta_{\text{start}}-\beta_{\text{end}})$, and $\sigma^{\prime}_{t}=\mu\sigma_{t},\mu=0.15$. While the standard self-diffusion loss relies solely on data fidelity, the joint estimation of $\mathbf{x}$ and $\mathbf{k}$ is a severely ill-posed problem that requires additional constraints. Therefore, we augment the loss with an $\ell_{1}$ term for the kernel. The networks are jointly optimized within an inner loop by minimizing the following objective:
$\mathcal{L}_{t}(\theta,\phi)={\|(D_{\theta}(\hat{\mathbf{x}}_{t})\circledast G_{\phi}(\hat{\mathbf{z}}_{t}))-\mathbf{y}\|_{2}^{2}}+{\lambda_{k}{R}(G_{\phi}(\hat{\mathbf{z}}_{t}))}$
@@ 3.2 Joint Image and Kernel Estimation
Algorithm 1 Blind Deblurring using self-diffusion (DeblurSDI)
@@
4:Kernel estimate $\mathbf{z}_{T}\sim\mathcal{N}(0,\mathbf{I})$; $G_{\phi}$ with random weights $\phi$
@@
11:   Generate kernel: $\mathbf{k}\leftarrow G_{\phi}(\hat{\mathbf{z}}_{t})$
@@
17:  Update kernel latent code: $\mathbf{z}_{t-1}\leftarrow G_{\phi}(\hat{\mathbf{z}}_{t})$
@@
19:return Reconstructed image $\mathbf{x}\leftarrow\mathbf{x}_{0}$ and final blur kernel $\mathbf{k}\leftarrow\mathbf{k}_{0}$
4 Implementation and Analysis

### v2

# Self-Diffusion Driven Blind Imaging

@@ Abstract
Optical imaging systems are inherently imperfect due to diffraction limits, lens manufacturing tolerances, assembly misalignment, and other physical constraints. In addition, unavoidable camera shake and object motion further introduce non-ideal degradations during acquisition. These aberrations and motion-induced variations are typically unknown, difficult to measure, and costly to model or calibrate in practice. Blind inverse problems offer a promising direction by jointly estimating both the latent image and the unknown degradation kernel. However, existing approaches often suffer from convergence instability, limited prior expressiveness, and sensitivity to hyperparameters. Inspired by recent advances in self-diffusion, we propose DeblurSDI, a zero-shot, self-supervised blind imaging framework that requires no pre-training. DeblurSDI formulates blind image recovery as an iterative reverse self-diffusion process that begins from pure noise and progressively refines both the sharp image and the blur kernel. Extensive experiments on combined optical aberrations and motion blur demonstrate that DeblurSDI consistently outperforms other methods by a substantial margin.
@@ 1 Introduction
Optical imaging systems are widely used in various sensors such as cameras [14, 6], microscopes [33], telescopes [30], and LiDARs that operate across wavelengths from visible to infrared. The lens system and the focal-plane image sensor jointly constitute a complete imaging module. In a paraxial, shift-invariant optical system, the captured image can be modeled as
$I(x,y)=O(x,y)\circledast h(x,y),$ (1)
where $O(x,y)$ denotes the object’s intensity distribution and $h(x,y)$ is the lens’s point spread function (PSF) given by
$h(x,y)=\left|\mathcal{F}\{P(\rho,\theta)\,e^{i2\pi W(\rho,\theta)/\lambda}\}\right|^{2}.$ (2)
Here $\mathcal{F}$ denotes the Fraunhofer diffraction operator [13], $P(\rho,\theta)$ is the complex pupil function, $W(\rho,\theta)$ is the wavefront, and $\lambda$ is the wavelength [5]. Because of inherent imperfections such as diffraction limits [13], manufacturing tolerances [34], assembly misalignment [26], and unavoidable camera shake and object motion [39, 15], the captured images are often corrupted by optical aberrations [26, 27] and motion blur [39, 35].
A straightforward way to improve image quality is to explicitly calibrate the system’s PSF by measuring optical aberrations and motion-blur kernels for each lens [30, 29, 33]. However, this calibration process is often time-consuming, requires dedicated hardware, and relies heavily on expert knowledge. Blind inverse problems offer a promising direction by jointly estimating both the latent image and the unknown degradation PSF [23, 10]. This approach provides a unified framework for both calibration and deblurring, and has been applied to a wide range of applications not limited to deblurring, such as super-resolution and image denoising.
However, existing approaches often exhibit unstable convergence [23], limited prior expressiveness [20], and pronounced sensitivity to hyperparameters [7, 28, 2, 21, 31], which make them difficult to use reliably in practice. Motivated by recent developments in self-diffusion [24], we propose DeblurSDI, a training-free and hyperparameter-insensitive blind imaging framework that achieves high stability while requiring no external priors. Inheriting the core idea of self-diffusion, which progressively restores images across hierarchical noise levels, DeblurSDI extends this principle to jointly recover the sharp image and PSF. This formulation significantly improves robustness in the otherwise unstable joint optimization, particularly in PSF estimation, which is the primary challenge in blind imaging problems.
Extensive experiments on simulated optical aberrations and motion blur demonstrate that DeblurSDI consistently outperforms state-of-the-art blind deconvolution methods by a substantial margin. Our conditions are summarized as follows: 1) We propose DeblurSDI, a training-free and self-supervised framework for blind image recovery that integrates self-diffusion and jointly estimates both the sharp image and the PSF; 2) We construct evaluation datasets that include optical-aberration PSFs and motion-blur kernels, enabling comprehensive assessment of blind imaging performance under challenging conditions; 3) We demonstrate that DeblurSDI achieves consistently superior performance and robustness compared to state-of-the-art blind deblurring methods, particularly in stability and PSF estimation.
@@ 2 Related Work
Correcting optical aberrations and motion blur generally follows one of two paradigms. In the model-based paradigm, the PSF is described using a parametric wavefront model such as Zernike polynomials [27, 25], and calibration procedures are employed to estimate the corresponding coefficients. This yields an explicit PSF that enables non-blind deconvolution. In the model-free paradigm, blind deconvolution is performed, which bypasses calibration entirely by estimating both the image and the PSF from the observation [1].
Optical Aberration Modeling and PSF Calibration:
Optical aberrations arise from diffraction limits [13], manufacturing tolerances [34], lens-group misalignment [26], and environmental variations [5], and they fundamentally shape the PSF of an imaging system. Early descriptions of aberrations were grounded in classical geometric optics [37], characterizing low-order distortions such as spherical aberration, coma, astigmatism, field curvature, and distortion [26]. With the development of wave optics [13], the PSF became formally modeled as the squared magnitude of the Fourier transform of the pupil function [16], where phase distortions encode the underlying wavefront errors. Modern optical systems typically represent these wavefront errors using Zernike polynomials [1, 27, 25], an orthonormal basis on the unit disk that directly corresponds to physically interpretable aberration modes and has become the industry standard for lens characterization [1].
Estimating the actual PSF of a physical lens, however, requires calibration procedures [30, 29]. In controlled laboratory settings, wavefront sensors such as Shack–Hartmann devices [30] provide direct measurements of Zernike coefficients, while phase-diversity methods infer aberrations from pairs of focused and defocused images [29, 12]. Microscopy systems often rely on submicron fluorescent beads [33] to sample spatially varying PSFs and construct eigen-PSF [32] or product-convolution models [3] for wide-field correction. Although accurate, these calibration pipelines require specialized hardware, multiple acquisitions, or carefully prepared samples, making them impractical for general-purpose cameras and consumer-level imaging devices.
As a result, industrial imaging systems such as smartphone cameras [6, 14], surveillance sensors, and automotive optical modules—typically rely on one-time factory calibration combined with lookup tables [6, 40] or compact neural modules embedded in the ISP pipeline. These approaches correct only approximate or averaged aberrations and cannot account for camera shake, object motion, or per-instance PSF variations in real scenes. Therefore, while model-based PSF calibration offers high fidelity, its cost, hardware dependency, and lack of adaptability highlight the need for calibration-free alternatives such as blind deconvolution methods [23, 10].
Blind Deconvolution with Training-free or Trained Priors:
Blind deconvolution aims to recover both the latent sharp image and the unknown blur kernel from a single degraded observation. Classical approaches typically formulate this as an alternating optimization problem that minimizes an energy functional over the image and the PSF [23, 10]. Early methods rely on handcrafted priors such as sparsity, heavy-tailed gradient distributions, or total variation to regularize the ill-posed joint estimation [20, 7]. Although effective in limited settings, these methods often exhibit strong sensitivity to initialization, kernel size, and hyperparameter choices, which leads to kernel drift and unstable convergence behaviors.
To alleviate the limitations of handcrafted priors, implicit neural priors have emerged as a powerful alternative. Deep Image Prior (DIP) demonstrates that the structure of an untrained convolutional network can serve as a natural image prior, enabling image restoration without external data [36]. Subsequent works extend this idea to blind deblurring by optimizing both the image and the kernel through coupled implicit networks [31]. These methods offer improved flexibility and avoid the need for large training datasets, yet the coupled optimization is still highly unstable and prone to overfitting, especially when the blur kernel is large or spatially complex.
Another line of work introduces pretrained generative models such as GANs, VAEs, and more recently diffusion models as strong image priors for inverse problems [4, 11, 38]. Diffusion-based solvers, including diffusion posterior sampling (DPS) and related score-based formulations, have demonstrated impressive performance in non-blind deblurring and other linear inverse problems [9, 18]. However, these methods rely heavily on large-scale pre-training and are restricted by domain mismatch between the training data and the target scene. More importantly, pretrained-prior approaches are generally incapable of estimating the blur kernel, and therefore cannot be directly applied to fully blind deconvolution.
Because existing methods either require dedicated calibration hardware or struggle with the inherent instability of joint image–kernel estimation. These limitations highlight the need for a framework that can both capture realistic optical degradations and robustly estimate the PSF without external training or calibration. Motivated by this gap, we now introduce our degradation modeling pipeline and the proposed DeblurSDI framework.
3 Method
In this section, we first introduce our aberration simulation framework, which generates a family of physically realistic PSFs derived from wavefront distortions modeled by Zernike polynomials. These simulated PSFs allow us to systematically evaluate blind imaging performance under controlled optical conditions. We then present DeblurSDI, a training-free blind deblurring framework that jointly recover the latent sharp image and the unknown PSF. By combining physically grounded PSF synthesis with a robust blind inverse solver, our approach enables comprehensive analysis and effective restoration across a wide range of optical aberrations and motion blur scenarios.
3.1 Modeling Optical Aberrations
To simulate realistic optical degradations, we adopt the standard wavefront–pupil–PSF formulation widely used in computational optics. Optical aberrations are represented through a weighted sum of Zernike polynomials [25], which form an orthonormal basis on the unit disk and are known to accurately model low- and high-order aberration modes such as defocus, astigmatism, coma, spherical aberration, trefoil, and quadrafoil.
Zernike-based wavefront modeling:
Let $(\rho,\theta)$ denote polar coordinates on the normalized pupil domain $\{(x,y):x^{2}+y^{2}\leq 1\}$. The Zernike polynomial of radial order $n$ and azimuthal order $m$ is defined as
[…]
where the radial term $R_{n}^{m}(\rho)$ is given by
$R_{n}^{m}(\rho)=\sum_{k=0}^{\frac{n-m}{2}}(-1)^{k}\,\frac{(n-k)!}{k!\,\big(\tfrac{n+m}{2}-k\big)!\,\big(\tfrac{n-m}{2}-k\big)!}\,\rho^{\,n-2k}.$ (4)
A wavefront corrupted by multiple aberrations can therefore be written as
$W(\rho,\theta)=\sum_{(n,m)\in\mathcal{A}}a_{n,m}\,Z_{n}^{m}(\rho,\theta),$ (5)
where each coefficient $a_{n,m}$ controls the severity of the corresponding aberration mode. In our simulations, we include all Zernike modes up to order $n=4$ (defocus, astigmatism, coma, trefoil, spherical, quadrafoil), which are sufficient to reproduce a wide range of realistic aberration patterns observed in lens-based imaging systems.
Pupil function and PSF generation:
Given the aberrated wavefront $W$, the complex pupil function is
$P(\rho,\theta)=\mathbf{1}_{\{\rho\leq 1\}}\,\exp\!\left(\frac{2\pi i}{\lambda}\,W(\rho,\theta)\right),$ (6)
where $\lambda$ denotes the wavelength and the indicator ensures that the phase modulation is applied only within the physical aperture. The PSF is obtained as the squared magnitude of the Fourier transform of the pupil:
$h(x,y)=\frac{\left|\mathcal{F}\!\left\{P(\rho,\theta)\right\}(x,y)\right|^{2}}{\max\limits_{x,y}\left|\mathcal{F}\!\left\{P(\rho,\theta)\right\}(x,y)\right|^{2}},$ (7)
where normalization ensures that $\max h=1$. We use an odd-sized Fourier grid so that the PSF peak is centered exactly at the middle pixel. A central crop of $h$ is used as the final convolution PSF in all experiments.
3.2 Blind Imaging via Self-Diffusion
@@
This section introduces the DeblurSDI method, which addresses blind optical correction by extending the self-diffusion framework [24]. The original self-diffusion method was designed for non-blind problems where the degradation operator is known. Our key contribution is to adapt this powerful framework to the more challenging blind imaging setting, where the degradation operator is also unknown. To achieve this, we introduce a novel process that jointly recovers the clean image and the PSF by optimizing two coupled, untrained neural networks within a self-contained reverse diffusion process.
Self-Diffusion:
@@
Self-diffusion is a pretraining free method designed to solve general linear inverse problems of the form $\mathcal{A}\mathbf{x}_{\text{true}}=\mathbf{y}$, where $\mathcal{A}$ is a known forward operator, $\mathbf{x}_{\text{true}}$ is the unknown solution, and $\mathbf{y}$ is the observation. It operates via an iterative reverse diffusion process that starts from pure Gaussian noise. At each step $t$, a noisy version of the current estimate $\mathbf{x}_{t}$ is created with
$\hat{\mathbf{x}}_{t}=\mathbf{x}_{t}+\sigma_{t}\cdot\epsilon_{t}~.$ (8)
@@
A single, randomly initialized network at the first time step, the self-denoiser $D_{\theta}$, is then optimized continuously by minimizing a data fidelity loss with respect to the original observation $\mathbf{y}$ for each time step $t$,
$\mathcal{L}_{t}(\theta)=\|\mathcal{A}D_{\theta,t}(\hat{\mathbf{x}}_{t})-\mathbf{y}\|_{2}^{2}~.$ (9)
@@
Joint Image and PSF Estimation:
Based on above observations, we apply such implicit regularization to PSF estimation. The forward model for blind image deblurring is given by
$\mathbf{y}=\mathbf{x}_{\text{true}}\circledast\mathbf{k}+n~,$ (10)
[…]
$\hat{\mathbf{x}}_{t}=\mathbf{x}_{t}+\sigma_{t}\cdot\epsilon_{x},\quad\text{and}\quad\hat{\mathbf{z}}_{t}=\mathbf{z}_{t}+\sigma^{\prime}_{t}\cdot\epsilon_{z},$ (11)
where $\epsilon_{x}\sim\mathcal{N}(0,\mathbf{I})$ and $\epsilon_{z}\sim\mathcal{N}(0,\mathbf{I})$. The noise schedule is $\sigma_{t}=\sqrt{1-\bar{\alpha}_{t}},~\text{where}~\bar{\alpha}_{t}=\mathop{\textstyle\prod}\nolimits_{i=0}^{t}(1-\beta_{i})~\text{and}~\beta_{t}=\beta_{\text{end}}+\frac{t}{T-1}(\beta_{\text{start}}-\beta_{\text{end}})$, and $\sigma^{\prime}_{t}=\mu\sigma_{t},\mu$ is a adjutable hyperparameter. It is often known that the PSF is to be sparse in practice. We thus define $R(\cdot)$ as L1-norm and use many layers of ReLU in the PSF generator to enforce sparsity on the generated PSF. The both networks are jointly optimized within an inner loop by minimizing the following objective:
$\mathcal{L}_{t}(\theta,\phi)={\|(D_{\theta}(\hat{\mathbf{x}}_{t})\circledast G_{\phi}(\hat{\mathbf{z}}_{t}))-\mathbf{y}\|_{2}^{2}}+{\lambda_{k}{R}(G_{\phi}(\hat{\mathbf{z}}_{t}))}$ (12)
@@
Algorithm 1 Blind Imaging via Self-Diffusion
@@
4:PSF estimate $\mathbf{z}_{T}\sim\mathcal{N}(0,\mathbf{I})$; $G_{\phi}$ with random weights $\phi$
@@
11:   Generate PSF: $\mathbf{k}\leftarrow G_{\phi}(\hat{\mathbf{z}}_{t})$
@@
17:  Update PSF latent code: $\mathbf{z}_{t-1}\leftarrow G_{\phi}(\hat{\mathbf{z}}_{t})$
@@
19:return Reconstructed image $\mathbf{x}\leftarrow\mathbf{x}_{0}$ and final PSF $\mathbf{k}\leftarrow\mathbf{k}_{0}$
4 Implementation and Experiments
4.1 Dataset
Optical Aberration:
The simulated optical aberration process is implemented numerically by evaluating the Zernike basis on a $255\times 255$ Cartesian grid, constructing the pupil (6), and computing the PSF via FFT according to (7). Figure 1 shows the resulting wavefronts and PSFs for individual and combined aberrations (defocus, coma, astigmatism, spherical, coma+spherical, defocus+coma). These aberrations are all applied on sharp images collected from above datasets to obtain blurred images. Then our proposed DeblurSDI is applied and evaluated on them.
Figure 1: PSFs for individual and combined aberrations (defocus, coma, astigmatism, spherical, coma+spherical, defocus+coma). Uprow: wavefronts. Downrow: PSFs.
Motion Blur:
To systematically evaluate the performance of our method, we collected four datasets including Levin [22], Cho [8], Kohler [19], and FFHQ [17]. The first three datasets are widely used benchmarks for blind image deblurring, containing various synthetic blur kernels applied to natural images. The FFHQ dataset consists of 20 random selected human face images and 4 blur kernels from Cho [8]. For our experiments, we use $T=30$ outer reverse diffusion steps. At each step, we run $S=200$ inner optimization iterations. For each dataset, we use the provided blur kernels and generate blurred images by convolving the sharp images with these kernels to simulate real-world conditions. Each blur kernel is applied to every image in the dataset. The details of our comprehensive datasets are shown in Table 1.
Table 1: Details of datasets for evaluation
Dataset Image Size Kernel Sizes
Levin1 $\times255$ {19, 17, 15, 27, 13, 21, 23, 23}
Cho2
[…]
{27, 23, 19, 21}
Kohler3 $\times800$
16, 14, 9, 13, 29, 17, 19, 98,
102, 62, 40, 29
FFHQ4 $\times256$ {27, 23, 19, 21}
Total Pairs 128